OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points (4, 5) and (–3, –1)? y - 1 = 6/7(x - 3) y + 1 = 6/7(x + 3) y + 3 = 7/6(x + 1) y + 1 = 6/5(x +3)

OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points (4, 5) and (–3, –1)?

OpenStudy (texaschic101):

first we find the slope by using the slope formula : slope (m) = (y2 - y1) / (x2 - x1) (4,5)...x1 = 4 and y1 = 5 (-3,-1)...x2 = -3 and y2 = -1 now sub slope(m) = (-1 - 5) / (-3 - 4) slope(m) = -6/-7 slope(m) = 6/7 now use point-slope formula y - y1 = m(x - x1) slope(m) = 6/7 use either of your points...(4,5)...x1 = 4 and y1 = 5 now sub y - 5 = 6/7(x - 4) OR slope(m) = 6/7 (-3,-1)....x1 = -3 and y1 = -1 y - (-1) = 6/7(x - (-3) y + 1 = 6/7(x + 3) <==== here it is

OpenStudy (anonymous):

thx